# Algorithms for Dynamic Computational Geometry with Applications

### v2

@@ Abstract
Most of the literature of computational geometry concerns geometric properties of sets of static points. M.J. Atallah introduced “dynamic computational geometry,” concerned with both momentary and long-term geometric properties of sets of moving point-objects. This area of research seems to have been dormant recently. The current paper examines new problems in dynamic computational geometry. Solutions are given for the sequential and coarse-grained multicomputer (CGM) models of computation.
@@ 1 Introduction
M.J. Atallah’s paper [1] presented efficient sequential algorithms for a variety of problems in computational geometry concerned with systems of point–objects in motion in Euclidean space ${\mathbb{R}}^{d}$; it was assumed for each object that its motion was described in each coordinate by known polynomials of time. Parallel versions of these algorithms were studied in [2, 3, 6, 7]. Another paper concerned with geometric properties of moving objects is [10]. The term “dynamic” is also applied to the study of a system of objects in which, over time, objects are inserted or deleted from the system, but this is not what is studied in the current paper.
In the current paper, we introduce efficient sequential and coarse grained parallel algorithms for additional problems in dynamic computational geometry (as introduced in Atallah’s paper [1]), and discuss applications.
@@ 2.2 On more general motion
Even if a library of functions containing the $\sin$ or $\cos$ is not available, the $\sin$ and $\cos$ functions can each be approximated by the sums of the first several terms of their respective appropriately centered Taylor series. Thus, motion as described in section (1) is included in the set of motions described in section 2.1. Similarly, a wide range of motion functions not described in the form of section 2.1 can be represented in this form with satisfactory approximation.
@@ 2.4 Coarse grained multicomputer
[4] Let $X=\{x_{i}\}_{i=1}^{n}$ be a set of elementary comparable data items distributed one per processor in a $CGM(n,p)$. The value of a semigroup operation (e.g.,  $\min\{x\in X\}$) can be computed and broadcast to all processors in optimal $\Theta(n/p)$ time.
@@ 3.1 Sequential solution
      Else If equation  $S_{j}(t)=d_{j}$   has a smallest solution $t_{j}\in[0,M]$ then
@@ 3.2 CGM solution
    Use Algorithm 3.1 so that each processor $P_{i}$ solves its portion of the problem, finding the earliest time $T_{i}$ that any of $\{q_{i}\}_{i=(j-1)n/p+1}^{jn/p}\}$ ($i\neq 1)$, if any, is too close to $q_{1}$.
@@
    Processor $P_{1}$ computes $m=\min\{T_{i}\}_{i=1}^{p}$ in $\Theta(p)$ time and notes the set $Y$ of indices at which the minimum occurs. I.e., $j\in Y$ if and only if $T_{j}=m$.
@@
This problem can be solved by algorithms similar to Algorithm 3.1 and 3.4. The sequential execution time is therefore a worst-case optimal $\Theta(n)$, and the execution time for a $CGM(n,p)$ is therefore a worst-case optimal $\Theta(n/p)$.
@@ 5.1 Preliminaries
  • Since we may have extended pieces to adjacent intervals, we may have undefined entries $(pieceFunction_{i},pieceInterval_{i})$. We can compress the defined entries to be consecutively indexed via a semigroup operation in $\Theta(s)=\Theta(1)$ time [9].
@@ 5.2 Solving the 3 aligned problem
Algorithm to find first (if any) instance of an $\varepsilon$-approximately collinear triple of $q_{1}$ and two other members of $Q$. Assume $\varepsilon>0$ is given.
@@
If $a(0)=0$ or $b(0)=0$ or $c(0)=0$ then tFirst = 0
@@ 6 Further remarks
We have studied the Too Close, Too Far, and 3 Aligned problems in dynamic computational geometry. We have given algorithms for these problems whose sequential running times are all linear in the amount of data to be processed. For the Too Close and Too Far problems, we have also given worst-case optimal solutions for the $CGM(n,p)$ model of parallel computation.

### v3

@@ Abstract
Most of the literature of computational geometry concerns geometric properties of sets of static points. M.J. Atallah introduced “dynamic computational geometry,” concerned with both momentary and long-term geometric properties of sets of moving point-objects. This area of research seems to have been dormant recently. The current paper examines new problems in dynamic computational geometry: the “Too Close,” “Too Far,” and “3 Aligned” problems. Worst-case optimal solutions are given for these problems using the sequential and coarse-grained multicomputer (CGM) models of computation.
@@ 1 Introduction
Computational geometry is generally concerned with geometric properties of some object that is usually a set $S$ of static points; less often, a set of line segments, curves, or surfaces. Examples of such properties include finding
  • the nearest pair of distinct members of $S$;
  • the diameter of $S$;
  • the convex hull of $S$;
  • maximal subsets of (exactly or approximately) collinear members of $S$
  • subsets of $S$ that are (exactly or approximately) congruent to, or similar to, a given set of points,
et al. Readers needing more background in computational geometry may find useful such sources as [12, 2, 9].
M.J. Atallah’s paper [1] introduced “dynamic computational geometry,” presenting efficient sequential algorithms for a variety of geometric problems concerned with systems of point–objects that are in simultaneous motion in Euclidean space ${\mathbb{R}}^{d}$; it was assumed for each object that its motion was described in each coordinate by known polynomials of time. Parallel versions of these algorithms were studied in [3, 4, 7, 8]. Another paper concerned with geometric properties of moving objects is [13]. The term “dynamic” is also applied to the study of a system of objects in which, over time, objects are inserted or deleted from the system, but this is not what is studied in the current paper.
In the current paper, we introduce efficient sequential and coarse grained parallel algorithms for additional problems in dynamic computational geometry (as introduced in [1]), and discuss applications. Problems considered are:
  • The “Too Close” problem: When, if ever, is the earliest time that one of our point objects, $q_{1}$, is too close to any other of our point objects, $q_{j}$? For moving objects, this seems to have safety implications.
  • The “Too Far” problem: When, if ever, is the earliest time that one of our point objects, $q_{1}$, is too far from any other of our point objects, $q_{j}$? This seems to have implications for the ability to communicate or the ability to attack.
  • The “3 Aligned” problem: When, if ever, is the earliest time that one of our point objects, $q_{1}$, is (approximately or exactly) collinear with any two others, $q_{i}$ and $q_{j}$. This seems to have implications for the ability to communicate or the ability to attack.
@@ 2.2 On more general motion
Even if a library of functions containing the $\sin$ or $\cos$ is not available, the $\sin$ and $\cos$ functions can each be approximated by the sums of the first several terms of their respective appropriately centered Taylor series. Thus, motion as described in (1) is included in the set of motions described in section 2.1. Similarly, a wide range of motion functions not described in the form of section 2.1 can be represented in this form with satisfactory approximation.
@@ 2.4 Coarse grained multicomputer
[5] Let $X=\{x_{i}\}_{i=1}^{n}$ be a set of elementary comparable data items distributed $n/p$ per processor in a $CGM(n,p)$. A parallel prefix operation on $X$ can be computed in optimal $\Theta(n/p)$ time.
@@ 3.1 Sequential solution
      Else If equation  $|S_{j}(t)|=d_{j}$   has a smallest solution $t_{j}\in[0,M]$ then
@@ 3.2 CGM solution
    Use Algorithm 3.1 so that each processor $P_{j}$ solves its portion of the problem, finding the earliest time $T_{j}$ that any of $\{q_{i}\}_{i=(j-1)n/p+1}^{jn/p}$ ($i\neq 1)$, if any, is too close to $q_{1}$.
@@
    Processor $P_{1}$ computes $m=\min\{T_{j}\}_{j=1}^{p}$ in $\Theta(p)$ time and notes the set $Y$ of indices at which the minimum occurs. I.e., $j\in Y$ if and only if $T_{j}=m$.
@@
This problem can be solved by algorithms similar to Algorithms 3.1 and 3.4. The sequential execution time is therefore a worst-case optimal $\Theta(n)$, and the execution time for a $CGM(n,p)$ is therefore a worst-case optimal $\Theta(n/p)$.
@@ 5.1 Preliminaries
  • Since we may have extended pieces to adjacent intervals, we may have undefined entries $(pieceFunction_{i},pieceInterval_{i})$. We can compress the defined entries to be consecutively indexed via a parallel prefix operation in $\Theta(s)=\Theta(1)$ time, by Theorem 2.7.
@@ 5.3 Sequential solution
Algorithm to find first (if any) instance of an $\varepsilon$-approximately collinear triple of $q_{1}$ and two other members of $Q$, sequentially. Assume $\varepsilon>0$ is given.
@@
If $|a(0)+b(0)-c(0)|\leq\varepsilon$ or $|a(0)+c(0)-b(0)|\leq\varepsilon$
or $|b(0)+c(0)-a(0)|\leq\varepsilon$ then tFirst = 0
@@ 6 Further remarks
We have studied the Too Close, Too Far, and 3 Aligned problems in dynamic computational geometry, for $n$ point-objects in motion as described in secction 2.1. We have given algorithms for these problems whose running times are all worst-case optimal. In particular, the running times are as follows.
$\begin{array}[]{lll}\underline{Problem}&\underline{Model~of~computation}&\underline{Worst-case~time}\\ Too~Close&Sequential&\Theta(n)\\ Too~Close&CGM(n,p)&\Theta(n/p)\\ Too~Far&Sequential&\Theta(n)\\ Too~Far&CGM(n,p)&\Theta(n/p)\\ 3~Aligned&Sequential&\Theta(n^{2})\end{array}$
We are grateful to the anonymous referees for useful suggestions.